OpenStudy (anonymous):

LOG3 (X-24) + LOG3 (X) = 4 I'm trying to solve the LOG problem above the I have 2 LOG problems I'm having difficulty with. With this one I have gotten down to 0= X^2 - 24X -81 but cant factor it, not sure if I've done it correctly :/ cuz I cant factor it

OpenStudy (jhannybean):

Remember: \(\log(a) + \log(b)=\log(ab)\)

OpenStudy (jhannybean):

\[a=(x-24) ~ , ~ b=x\]

OpenStudy (anonymous):

Yes, what I have so far is LOG2 X(X-24)=4 then LOG3 (X^2 -24X) =4 3^4 = X^2 -24X 0 = X^2 -24X - 81

OpenStudy (anonymous):

I think the next step is to factor it but I cant so I'm guessing I did something wrong

OpenStudy (jhannybean):

Alright, let me work it out, checking.

OpenStudy (jhannybean):

Ok, think of factors that will multiply to give you -81 but add to give you -24 hint: 3 is one of them

OpenStudy (anonymous):

O! +3 and -27!

OpenStudy (jhannybean):

:)

OpenStudy (anonymous):

Thank you!!

OpenStudy (jhannybean):

No problem, haha.